### v1

@@ 4.1.2 Low-level vision
In the realm of image super-resolution, Meet More Areas (MMAcheng2024activating stands out as a novel model designed for super-resolution tasks. Built on the ViM block, MMA aims to enhance performance by activating a wider range of areas within images. To achieve this, MMA adopts several key strategies, including integrating ViM into MetaFormer-style modules, pre-training ViM on larger datasets, and employing complementary attention mechanisms. MMA comprises three main modules: shallow feature extraction, deep feature extraction, and high-quality reconstruction. Leveraging the ViM module, MMA effectively models global information and further expands the activation region through attention mechanisms.
@@
Existing restoration backbones often face the dilemma between global receptive fields and efficient computation, hindering their application in practice, while Mamba shows great potential for long-range dependency modeling with linear complexity, which offers a way to resolve the above dilemma. MambaIR guo2024mambair aims to solve the problem by introducing local enhancement and channel attention mechanisms to improve the standard Mamba model. The methodology of the model mainly consists of three stages: shallow feature extraction, deep feature extraction, and high-quality image reconstruction. Among them, the deep feature extraction stage utilizes multiple residual state space blocks (RSSBs) for feature extraction, which add a VSS block before the channel attention block designed in previous transformer-based restoration networks. SERPENT shahab2024serpent designs a hierarchical architecture that processes input images in a multi-scale manner, including processing steps such as segmentation, embedding, downsampling, and upsampling, and introduces jump connections to facilitate information flow. Among them, the Serpent block is the main processing unit, consisting of multiple VSS blocks stacked on each other. Serpent combines the advantages of Convolutional Networks and Transformers to drastically reduce the computational effort, GPU memory requirement, and model size while maintaining high reconstruction quality. VmambaIR shi2024vmambair proposes the OSS module to comprehensively and efficiently model image features from six directions. The omnidirectional selective scanning mechanism overcomes the unidirectional modeling limitation of SSMs and achieves comprehensive pattern recognition and modeling by modeling the image information flow in all three dimensions.
@@
UVM-Net zheng2024u is a novel single-image defogging network architecture that combines the local feature extraction of convolutional layers and SSM’s long-range dependency modeling capability to exhibit efficient performance. The method employs an encoder-decoder network architecture, and the critical component is the ViM block, which leverages the long-range modeling capability of SSM by rolling the feature map over the channel domain. Unlike U-Mamba ma2024u and Mamba-UNet  wang2024mamba, the ViM block establishes long-range dependencies on another dimension of the feature map (the non-channel domain).
@@
Images lose important frequency information under the influence of raindrops, which affects the performance of visual perception and advanced visual tasks.FreqMamba zhen2024freqmamba is a novel image de-raining method that combines Mamba modeling and frequency analysis techniques to solve the image de-raining problem. Specifically, FreqMamba contains three branching structures, including spatial Mamba, frequency band Mamba, and Fourier global modeling. Spatial Mamba processes raw image features to extract details and correlations within the image. Frequency Band Mamba uses the Wavelet Packet Transform (WPT) to decompose the input features into spectral features in different frequency bands and scan them along the frequency dimension. Fourier modeling i.e., processing the input using the Fourier transform captures the global degradation patterns affecting the image. Extensive experiments have shown that FreqMamba outperforms existing state-of-the-art methods both visually and quantitatively.
@@
Image deblurring is a classical problem in low-level computer vision, which aims to recover high-quality clear images from blurred input images.ALGNet gao2024aggregating is an efficient image deblurring network that utilizes selective state-space models (SSM) to aggregate rich and accurate features.The network consists of multiple ALGBlocks, each of which contains a CLGF module that captures local and global features and a feature aggregation module FA.The CLGF module captures long-range dependent features using SSM and employs a channel-attention mechanism to reduce local pixel forgetting and channel redundancy.The FA module emphasizes the importance of local features in recovery by recalibrating the weights.
@@
The efficiency of Mamba contributes significantly to mitigating the high computational complexity associated with training generation tasks. To address the change in generating long and diverse sequences with low latency, MambaTalk xu2024mambatalk implements a two-stage modeling strategy with discrete motion priors to enhance the quality of gestures and employs mamba block to enhance gesture diversity and rhythm through multimodal integration. Motion Mamba zhang2024motion is introduced to construct a motion generation model based on Mamba, leveraging its efficient hardware-aware design. Motion Mamba consists of two main components: the Hierarchical Temporal Mamba (HTM) block for temporal data handling, and the Bidirectional Spatial Mamba (BSM) block for processing latent poses. The HTM block employs several isolated SSM modules within a symmetric U-Net architecture to maintain motion consistency across frames. Meanwhile, the BSM block enhances the accuracy of motion generation within a temporal frame by processing latent poses bidirectionally. Diffusion State Space Models (DiS)  fei2024scalable substitute the conventional U-Net backbone in diffusion models with SSMs. This framework considers all inputs, including time, conditions, and noisy image patches and tokens. To tackle the oversight of spatial continuity in the scanning scheme of existing Mamba-based vision methods, Zigzag Mamba hu2024zigma is introduced as a straightforward, plug-and-play solution inspired by DiT style approaches. Essentially, it retains the scanning scheme of Plain Mamba but expands it from four to eight schemes by incorporating their mirror flipping schemes, as shown in (f) Fig. 4. Subsequently, Zigzag Mamba is integrated with the Stochastic Interpolant framework, forming ZigMa, to explore the scalability of the diffusion model on large-resolution visual datasets.GAMBA shen2024gamba introduces a sequential network based on Mamba, which enables context-dependent reasoning and linear scalability for sequence length. This architecture accommodates many Gaussians for the 3D Gaussian splatting process. To address the issue of quadratic memory consumption increase with sequence length in traditional attention-based video generative diffusion models, SSM-based diffusion model oshima2024ssm is introduced for generating longer video sequences. Like ViS4mer  islam2022long, the SSM-based diffusion model reimagines the attention modules within the conventional temporal layers of Video Diffusion Models (VDMs). It replaces them with a ViM block designed to capture the temporal dynamics of video data, accompanied by a Multi-Layer Perceptron (MLP) to boost model performance. This innovative approach significantly mitigates memory consumption for extended sequences.
@@
The irregularity and sparsity of point cloud data have been a challenge in 3D vision. Although Transformer shows potential in point cloud analysis tasks based on its powerful global information modeling capability, its computational complexity grows significantly with the increase of input length, limiting its application to long sequence models. In this context, SSPoint Mamba liang2024pointmamba is proposed as a simple and effective state space model for point cloud analysis. The model uses embedded point blocks as inputs and enhances the global modeling capability of SSM with a reordering strategy that provides a more rational geometric scanning order. The reordered point tokens (point tokens) are fed into a series of Mamba blocks to causally capture the point cloud structure. The model demonstrates its effectiveness on several point cloud analysis tasks.3DMambaComplete li20243dmambacomplete network aims to address the computational complexity challenges posed by the loss of local details and attention mechanisms common in point cloud completion, and it is built on top of the novel Mamba framework. The method first downsamples incomplete point clouds, then enhances feature learning with Mamba Encoder, then predicts and refines hyperpoints, and disperses hyperpoints to different 3D locations by learning specific offsets, and finally performs point deformation to generate a complete point cloud. The model utilizes the concept of structured state-space modeling to improve shape reconstruction by predicting hyperpoints and point deformation to control the deformation at each hyperpoint location.The point cloud filtering model 3DMambaIPF zhou20243dmambaipf, on the other hand, aims to deal with the denoising of large-scale point cloud data. The network integrates Mamba into a filtering module, Mamba-Denoise, to achieve accurate and fast modeling of long sequences of point cloud features.3DMambaIPF consists of several Mamba-Denoise modules, and employs a strategy of iterative point cloud filtering to implement the filtering process for point clouds. The loss function includes reconstruction loss and differentiable rendering loss to minimize the distance between the noisy point cloud and the real point cloud, optimize the visual boundary of the point cloud, and improve the realism of the denoising results.Point Cloud Mamba zhang2024point combines local and global modeling frameworks and proposes a Consistent Traversal Serialization (CTS) approach to transform 3D point cloud data into 1D point sequences while ensuring that adjacent points in the sequence are also spatially adjacent. In addition, point cueing and position encoding based on spatial coordinate mapping are introduced to help Mamba process point sequences more efficiently and inject position information.Point Mamba liu2024point is a new backbone network for point cloud processing that addresses the causality of state space models on point cloud data by introducing an octree-based ordering strategy. In addition, Point Mamba blocks incorporate a bi-directional selective scanning mechanism to adjust the sequence order dependency of Mamba.
@@ 4.2 Medical Visual Mamba
Transformers vaswani2017attention have profoundly influenced the field of medical imaging with their ability to master complex data representations. They have led to notable advancements across various imaging modalities, including Radiography seeram2019digital, Endoscopy lui2020overview, Computed Tomography (CT) withers2021x, Ultrasound Images christensen2020super, and Magnetic Resonance Imaging (MRI) tiwari2020brain. However, because most medical images are high-resolution and detailed, transformer models typically require considerable computational resources, which scale quadratically with image resolution.
@@
-2cm0cm Category Sub-category Method Efficiency Code 2D Segmentation Mamba-UNet  wang2024mamba - H-vmunet  wu2024h Memory 0.676 Params 8.97 Mamba-HUNet  shahriar2024integrating - P-Mamba  ye2024p Inference speed 23.49 Memory 12.22 Params 183.37 FLOPs 71.81×$10^{9}$ ProMamba  xie2024promamba Params 102 TM-UNet  tang2024rotate Params 14.86 Total Params 8.41 FLOPs 3.42 Semi-Mamba-UNet  ma2024semi - Swin-UMamba  liu2024swin Params 28 FLOPs 18.9 UltraLight VM-UNet  wu2024ultralight Params 0.049 GFLOPs 0.060 U-Mamba  ma2024u - VM-UNet  ruan2024vm Params 34.62 FLOPs 7.56 FPS 20.612 VM-UNET-V2  zhang2024vm Params 17.91 FLOPS 4.40 FPS 32.58 Weak-Mamba-UNet  wang2024weak - Radiation dose prediction MD-Dose  fu2024md Inference speed 18 Params 30.47 Classification MedMamba  yue2024medmamba - MambaMIL  yang2024mambamil - Image reconstruction MambaMIR/MambaMIR-GAN  huang2024mambamir - Exposure correction FDVM-Net  zheng2024fd Inference speed 22.95 3D Segmentation LMa-UNet  wang2024large - LightM-UNet  liao2024lightm Params 1.87 FLOPs 457.62×$10^{9}$ SegMamba  xing2024segmamba Inference speed 151 T-Mamba  hao2024t - Vivim  yang2024vivim FPS 35.33 Classification CMViM yang2024cmvim Params 50 Motion tracking Motion-Guided Dual-Camera Tracker  zhang2024motion2 - Backbone nnMamba  gong2024nnmamba Params 15.55 FLOPs 141.14 Image registration VMambaMorph  wang2024vmambamorph Inference speed 19 Memory 3.93 Params 9.64 MambaMorph  guo2024mambamorph Inference speed 27 Memory 7.60 Params 7.59 {tablenotes}
@@
Most of the innovative architectures that have been developed are based on U-Net, which has achieved remarkable success in various medical image segmentation tasks. U-Mamba  ma2024u is the first extension of the mamba model to the U-Net architecture for visual segmentation in biomedical imaging, addressing the challenge of long-range dependency modeling, which is captured by a hybrid CNN-SSM block. Wu et al. introduced the High-order Vision Mamba UNet (H-vmunet)  wu2024h, an improvement of U-Mamba, which utilizes a High-order 2D-selective-scan at each interaction order to bolster the learning of local features while minimizing the incorporation of redundant information. Shortly after their initial release, the team expanded their work by introducing the UltraLight VM-UNet  wu2024ultralight. This new iteration was developed by in-depth analysis of the critical factors affecting parameter efficiency within the Mamba framework. This resulted in a remarkably lightweight model with a mere 0.049M parameters and a computational efficiency of only 0.060 GFLOPs.In addition, Mamba-UNet wang2024mamba combines the encoder-decoder architecture of U-Net with the capabilities of Mamba and maintains spatial information at different network scales through jump connections. A Visual Mamba-based VSS block is used, which utilizes linear embedding layers and deep convolution to extract features while downsampling and upsampling are facilitated by multiple merge operations and extension layers for comprehensive feature learning.
@@
Pyramid ViT (PVT) and Swin-Unet are both pioneering hierarchical designs that apply visual tasks and propose progressive shrinking pyramids and spatial-reduction attention. Drawing inspiration from PVT and Swin-Unet, Ruan et al. introduced VM-UNet  ruan2024vm, a foundational model for purely SSM-based segmentation in medical imaging. This model demonstrates the capabilities of SSM in medical image segmentation and consists of three primary components: an encoder, a decoder, and skip connections. Building on their previous work, the team proposed VM-UNET-V2  zhang2024vm. The Visual State Space (VSS) block was introduced to capture a broader range of contextual information. The Semantics and Detail Infusion (SDI) mechanism was also implemented to enhance the fusion of low-level and high-level features. Mamba-HUNet  shahriar2024integrating, another multi-scale hierarchical upsampling network, incorporates the Mamba technique. It leverages Visual State Space blocks and patch merging layers to extract hierarchical features, ensuring the preservation of spatial information. TM-UNet  tang2024rotate introduces improvements to the bottleneck layer. This architecture proposes Triplet SSM as the bottleneck layer, representing the first attempt to use a pure SSM approach to integrate spatial and channel features. Existing Mamba-based models are primarily trained from scratch, overlooking potential benefits. Swin-UMamba  liu2024swin, a new Mamba-based model tailored explicitly for medical image segmentation tasks, leveraged the strengths of ImageNet-based pretraining.
@@
Previous discussions have primarily focused on supervised learning methods, but other supervisory approaches have also been explored. Semi-Mamba-UNet  ma2024semi combines a visual Mamba-based U-Shape Encoder-Decoder with a traditional CNN-based UNet in a semi-supervised learning framework. It introduces a self-supervised pixel-level contrastive learning strategy using a pair of projectors to improve feature learning, particularly on unlabeled data. Weak-Mamba-UNet  wang2024weak is a novel weakly-supervised learning framework for medical image segmentation, combining CNNs, ViT, and the VMamba. It focuses on scribble-based annotations and uses a collaborative, cross-supervisory mechanism with pseudo labels for iterative network learning and refinement.
@@
Some segmentation approaches diverge from UNet architectures. P-Mamba  ye2024p introduces a novel dual-branch framework for highly efficient left ventricle segmentation in pediatric echocardiograms. This model features an innovative DWT-based encoder branch equipped with Perona-Malik Diffusion (PMD) Blocks. Moreover, P-Mamba adopts vision mamba layers within its vision mamba encoder branch to bolster computational and memory efficiency. PromptMamba  xie2024promamba represents a groundbreaking integration of the Vision-Mamba and prompt technologies, marking a significant milestone as the first model to leverage the Mamba framework for the specific task of polyp segmentation.
@@
In addition, mamba has expanded its research in 2D medical imaging beyond segmentation, enhancing the precision and speed of image analysis to support diagnosis and treatment planning. Classification is a vital and fundamental task in the field of medical image analysis. Yue et al. introduced Vision Mamba for this purpose, also known as MedMamba  yue2024medmamba. They developed a novel module named Conv-SSM, which combines the local feature extraction of convolutional layers with the long-range dependency capture of SSM, enabling efficient modeling of medical images from different modalities. Furthermore, MambaMIL  yang2024mambamil introduces the Sequence Reordering Mamba (SR-Mamba), a model that recognizes the order and distribution of instances in long sequences to harness the embedded valuable information effectively. Image reconstruction is pivotal in enhancing diagnostic processes, as high-quality and high-fidelity medical images are fundamental to the accuracy and efficiency of clinical decisions. Huang et al.  huang2024mambamir have developed MambaMIR, a model leveraging Mamba technology for the reconstruction of medical images, alongside its advanced counterpart, MambaMIR-GAN, which incorporates Generative Adversarial Networks. Moreover, Zheng et al. introduce FDVision Mamba (FDVM-Net)  zheng2024fd, a frequency-domain-based network that effectively corrects image exposure by reconstructing the frequency domain of endoscopic images, as recorded endoscopic images often suffer from exposure abnormalities. In specialized areas, MD-Dose  fu2024md, a cutting-edge diffusion model based on the Mamba architecture, was designed to predict radiation therapy dose distribution for thoracic cancer patients accurately.
@@
3D image analysis in medical imaging enables more accurate and comprehensive diagnoses by providing detailed views of complex anatomical structures. Gong et al. present nnMamba  gong2024nnmamba, an innovative architecture designed for 3D medical imaging applications, which integrates local and global relationship modeling via the MICCSS (Mamba-In-Convolution with Channel-Spatial Siamese input) module. nnMamba was tested on a comprehensive benchmark of six datasets for three crucial tasks: segmentation, classification, and landmark detection, showcasing its capability in long-range relationship modeling at channel and spatial levels.

### v2

@@ 4.1.2 Low-level vision
In the realm of image super-resolution, Meet More Areas (MMA[41] stands out as a novel model designed for super-resolution tasks. Built on the ViM block, MMA aims to enhance performance by activating a wider range of areas within images. To achieve this, MMA adopts several key strategies, including integrating ViM into MetaFormer-style modules, pre-training ViM on larger datasets, and employing complementary attention mechanisms. MMA comprises three main modules: shallow feature extraction, deep feature extraction, and high-quality reconstruction. Leveraging the ViM module, MMA effectively models global information and further expands the activation region through attention mechanisms.
@@
Existing restoration backbones often face the dilemma between global receptive fields and efficient computation, hindering their application in practice, while Mamba shows great potential for long-range dependency modeling with linear complexity, which offers a way to resolve the above dilemma. MambaIR [57] aims to solve the problem by introducing local enhancement and channel attention mechanisms to improve the standard Mamba model. The methodology of the model mainly consists of three stages: shallow feature extraction, deep feature extraction, and high-quality image reconstruction. Among them, the deep feature extraction stage utilizes multiple residual state space blocks (RSSBs) for feature extraction, which add a VSS block before the channel attention block designed in previous transformer-based restoration networks. SERPENT [58] designs a hierarchical architecture that processes input images in a multi-scale manner, including processing steps such as segmentation, embedding, downsampling, and upsampling, and introduces jump connections to facilitate information flow. Among them, the Serpent block is the main processing unit, consisting of multiple VSS blocks stacked on each other. Serpent combines the advantages of Convolutional Networks and Transformers to drastically reduce the computational effort, GPU memory requirement, and model size while maintaining high reconstruction quality. VmambaIR [35] proposes the OSS module to comprehensively and efficiently model image features from six directions. The omnidirectional selective scanning mechanism overcomes the unidirectional modeling limitation of SSMs and achieves comprehensive pattern recognition and modeling by modeling the image information flow in all three dimensions.
@@
UVM-Net [75] is a novel single-image defogging network architecture that combines the local feature extraction of convolutional layers and SSM’s long-range dependency modeling capability to exhibit efficient performance. The method employs an encoder-decoder network architecture, and the critical component is the ViM block, which leverages the long-range modeling capability of SSM by rolling the feature map over the channel domain. Unlike U-Mamba [85] and Mamba-UNet  [51], the ViM block establishes long-range dependencies on another dimension of the feature map (the non-channel domain).
@@
Images lose important frequency information under the influence of raindrops, which affects the performance of visual perception and advanced visual tasks.FreqMamba [68] is a novel image de-raining method that combines Mamba modeling and frequency analysis techniques to solve the image de-raining problem. Specifically, FreqMamba contains three branching structures, including spatial Mamba, frequency band Mamba, and Fourier global modeling. Spatial Mamba processes raw image features to extract details and correlations within the image. Frequency Band Mamba uses the Wavelet Packet Transform (WPT) to decompose the input features into spectral features in different frequency bands and scan them along the frequency dimension. Fourier modeling i.e., processing the input using the Fourier transform captures the global degradation patterns affecting the image. Extensive experiments have shown that FreqMamba outperforms existing state-of-the-art methods both visually and quantitatively.
@@
Image deblurring is a classical problem in low-level computer vision, which aims to recover high-quality clear images from blurred input images.ALGNet [76] is an efficient image deblurring network that utilizes selective state-space models (SSM) to aggregate rich and accurate features.The network consists of multiple ALGBlocks, each of which contains a CLGF module that captures local and global features and a feature aggregation module FA.The CLGF module captures long-range dependent features using SSM and employs a channel-attention mechanism to reduce local pixel forgetting and channel redundancy.The FA module emphasizes the importance of local features in recovery by recalibrating the weights.
@@
The efficiency of Mamba contributes significantly to mitigating the high computational complexity associated with training generation tasks. To address the change in generating long and diverse sequences with low latency, MambaTalk [77] implements a two-stage modeling strategy with discrete motion priors to enhance the quality of gestures and employs mamba block to enhance gesture diversity and rhythm through multimodal integration. Motion Mamba [37] is introduced to construct a motion generation model based on Mamba, leveraging its efficient hardware-aware design. Motion Mamba consists of two main components: the Hierarchical Temporal Mamba (HTM) block for temporal data handling, and the Bidirectional Spatial Mamba (BSM) block for processing latent poses. The HTM block employs several isolated SSM modules within a symmetric U-Net architecture to maintain motion consistency across frames. Meanwhile, the BSM block enhances the accuracy of motion generation within a temporal frame by processing latent poses bidirectionally. Diffusion State Space Models (DiS)  [78] substitute the conventional U-Net backbone in diffusion models with SSMs. This framework considers all inputs, including time, conditions, and noisy image patches and tokens. To tackle the oversight of spatial continuity in the scanning scheme of existing Mamba-based vision methods, Zigzag Mamba [34] is introduced as a straightforward, plug-and-play solution inspired by DiT style approaches. Essentially, it retains the scanning scheme of Plain Mamba but expands it from four to eight schemes by incorporating their mirror flipping schemes, as shown in (f) Fig. 4. Subsequently, Zigzag Mamba is integrated with the Stochastic Interpolant framework, forming ZigMa, to explore the scalability of the diffusion model on large-resolution visual datasets.GAMBA [83] introduces a sequential network based on Mamba, which enables context-dependent reasoning and linear scalability for sequence length. This architecture accommodates many Gaussians for the 3D Gaussian splatting process. To address the issue of quadratic memory consumption increase with sequence length in traditional attention-based video generative diffusion models, SSM-based diffusion model [84] is introduced for generating longer video sequences. Like ViS4mer  [72], the SSM-based diffusion model reimagines the attention modules within the conventional temporal layers of Video Diffusion Models (VDMs). It replaces them with a ViM block designed to capture the temporal dynamics of video data, accompanied by a Multi-Layer Perceptron (MLP) to boost model performance. This innovative approach significantly mitigates memory consumption for extended sequences.
@@
The irregularity and sparsity of point cloud data have been a challenge in 3D vision. Although Transformer shows potential in point cloud analysis tasks based on its powerful global information modeling capability, its computational complexity grows significantly with the increase of input length, limiting its application to long sequence models. In this context, SSPoint Mamba [82] is proposed as a simple and effective state space model for point cloud analysis. The model uses embedded point blocks as inputs and enhances the global modeling capability of SSM with a reordering strategy that provides a more rational geometric scanning order. The reordered point tokens (point tokens) are fed into a series of Mamba blocks to causally capture the point cloud structure. The model demonstrates its effectiveness on several point cloud analysis tasks.3DMambaComplete [79] network aims to address the computational complexity challenges posed by the loss of local details and attention mechanisms common in point cloud completion, and it is built on top of the novel Mamba framework. The method first downsamples incomplete point clouds, then enhances feature learning with Mamba Encoder, then predicts and refines hyperpoints, and disperses hyperpoints to different 3D locations by learning specific offsets, and finally performs point deformation to generate a complete point cloud. The model utilizes the concept of structured state-space modeling to improve shape reconstruction by predicting hyperpoints and point deformation to control the deformation at each hyperpoint location.The point cloud filtering model 3DMambaIPF [80], on the other hand, aims to deal with the denoising of large-scale point cloud data. The network integrates Mamba into a filtering module, Mamba-Denoise, to achieve accurate and fast modeling of long sequences of point cloud features.3DMambaIPF consists of several Mamba-Denoise modules, and employs a strategy of iterative point cloud filtering to implement the filtering process for point clouds. The loss function includes reconstruction loss and differentiable rendering loss to minimize the distance between the noisy point cloud and the real point cloud, optimize the visual boundary of the point cloud, and improve the realism of the denoising results.Point Cloud Mamba [81] combines local and global modeling frameworks and proposes a Consistent Traversal Serialization (CTS) approach to transform 3D point cloud data into 1D point sequences while ensuring that adjacent points in the sequence are also spatially adjacent. In addition, point cueing and position encoding based on spatial coordinate mapping are introduced to help Mamba process point sequences more efficiently and inject position information.Point Mamba [44] is a new backbone network for point cloud processing that addresses the causality of state space models on point cloud data by introducing an octree-based ordering strategy. In addition, Point Mamba blocks incorporate a bi-directional selective scanning mechanism to adjust the sequence order dependency of Mamba.
@@ 4.2 Medical Visual Mamba
Transformers [8] have profoundly influenced the field of medical imaging with their ability to master complex data representations. They have led to notable advancements across various imaging modalities, including Radiography [86], Endoscopy [87], Computed Tomography (CT) [88], Ultrasound Images [89], and Magnetic Resonance Imaging (MRI) [90]. However, because most medical images are high-resolution and detailed, transformer models typically require considerable computational resources, which scale quadratically with image resolution.
@@
-2cm0cm Category Sub-category Method Efficiency Code 2D Segmentation Mamba-UNet  [51] - H-vmunet  [55] Memory 0.676 Params 8.97 Mamba-HUNet  [59] - P-Mamba  [91] Inference speed 23.49 Memory 12.22 Params 183.37 FLOPs 71.81×$10^{9}$ ProMamba  [92] Params 102 TM-UNet  [60] Params 14.86 Total Params 8.41 FLOPs 3.42 Semi-Mamba-UNet  [52] - Swin-UMamba  [61] Params 28 FLOPs 18.9 UltraLight VM-UNet  [62] Params 0.049 GFLOPs 0.060 U-Mamba  [85] - VM-UNet  [63] Params 34.62 FLOPs 7.56 FPS 20.612 VM-UNET-V2  [64] Params 17.91 FLOPS 4.40 FPS 32.58 Weak-Mamba-UNet  [93] - Radiation dose prediction MD-Dose  [94] Inference speed 18 Params 30.47 Classification MedMamba  [65] - MambaMIL  [95] - Image reconstruction MambaMIR/MambaMIR-GAN  [56] - Exposure correction FDVM-Net  [96] Inference speed 22.95 3D Segmentation LMa-UNet  [45] - LightM-UNet  [97] Params 1.87 FLOPs 457.62×$10^{9}$ SegMamba  [98] Inference speed 151 T-Mamba  [99] - Vivim  [38] FPS 35.33 Classification CMViM [100] Params 50 Motion tracking Motion-Guided Dual-Camera Tracker  [46] - Backbone nnMamba  [101] Params 15.55 FLOPs 141.14 Image registration VMambaMorph  [53] Inference speed 19 Memory 3.93 Params 9.64 MambaMorph  [102] Inference speed 27 Memory 7.60 Params 7.59
  • 1
@@
Most of the innovative architectures that have been developed are based on U-Net, which has achieved remarkable success in various medical image segmentation tasks. U-Mamba  [85] is the first extension of the mamba model to the U-Net architecture for visual segmentation in biomedical imaging, addressing the challenge of long-range dependency modeling, which is captured by a hybrid CNN-SSM block. Wu et al. introduced the High-order Vision Mamba UNet (H-vmunet)  [55], an improvement of U-Mamba, which utilizes a High-order 2D-selective-scan at each interaction order to bolster the learning of local features while minimizing the incorporation of redundant information. Shortly after their initial release, the team expanded their work by introducing the UltraLight VM-UNet  [62]. This new iteration was developed by in-depth analysis of the critical factors affecting parameter efficiency within the Mamba framework. This resulted in a remarkably lightweight model with a mere 0.049M parameters and a computational efficiency of only 0.060 GFLOPs.In addition, Mamba-UNet [51] combines the encoder-decoder architecture of U-Net with the capabilities of Mamba and maintains spatial information at different network scales through jump connections. A Visual Mamba-based VSS block is used, which utilizes linear embedding layers and deep convolution to extract features while downsampling and upsampling are facilitated by multiple merge operations and extension layers for comprehensive feature learning.
@@
Pyramid ViT (PVT) and Swin-Unet are both pioneering hierarchical designs that apply visual tasks and propose progressive shrinking pyramids and spatial-reduction attention. Drawing inspiration from PVT and Swin-Unet, Ruan et al. introduced VM-UNet  [63], a foundational model for purely SSM-based segmentation in medical imaging. This model demonstrates the capabilities of SSM in medical image segmentation and consists of three primary components: an encoder, a decoder, and skip connections. Building on their previous work, the team proposed VM-UNET-V2  [64]. The Visual State Space (VSS) block was introduced to capture a broader range of contextual information. The Semantics and Detail Infusion (SDI) mechanism was also implemented to enhance the fusion of low-level and high-level features. Mamba-HUNet  [59], another multi-scale hierarchical upsampling network, incorporates the Mamba technique. It leverages Visual State Space blocks and patch merging layers to extract hierarchical features, ensuring the preservation of spatial information. TM-UNet  [60] introduces improvements to the bottleneck layer. This architecture proposes Triplet SSM as the bottleneck layer, representing the first attempt to use a pure SSM approach to integrate spatial and channel features. Existing Mamba-based models are primarily trained from scratch, overlooking potential benefits. Swin-UMamba  [61], a new Mamba-based model tailored explicitly for medical image segmentation tasks, leveraged the strengths of ImageNet-based pretraining.
@@
Previous discussions have primarily focused on supervised learning methods, but other supervisory approaches have also been explored. Semi-Mamba-UNet  [52] combines a visual Mamba-based U-Shape Encoder-Decoder with a traditional CNN-based UNet in a semi-supervised learning framework. It introduces a self-supervised pixel-level contrastive learning strategy using a pair of projectors to improve feature learning, particularly on unlabeled data. Weak-Mamba-UNet  [93] is a novel weakly-supervised learning framework for medical image segmentation, combining CNNs, ViT, and the VMamba. It focuses on scribble-based annotations and uses a collaborative, cross-supervisory mechanism with pseudo labels for iterative network learning and refinement.
@@
Some segmentation approaches diverge from UNet architectures. P-Mamba  [91] introduces a novel dual-branch framework for highly efficient left ventricle segmentation in pediatric echocardiograms. This model features an innovative DWT-based encoder branch equipped with Perona-Malik Diffusion (PMD) Blocks. Moreover, P-Mamba adopts vision mamba layers within its vision mamba encoder branch to bolster computational and memory efficiency. PromptMamba  [92] represents a groundbreaking integration of the Vision-Mamba and prompt technologies, marking a significant milestone as the first model to leverage the Mamba framework for the specific task of polyp segmentation.
@@
In addition, mamba has expanded its research in 2D medical imaging beyond segmentation, enhancing the precision and speed of image analysis to support diagnosis and treatment planning. Classification is a vital and fundamental task in the field of medical image analysis. Yue et al. introduced Vision Mamba for this purpose, also known as MedMamba  [65]. They developed a novel module named Conv-SSM, which combines the local feature extraction of convolutional layers with the long-range dependency capture of SSM, enabling efficient modeling of medical images from different modalities. Furthermore, MambaMIL  [95] introduces the Sequence Reordering Mamba (SR-Mamba), a model that recognizes the order and distribution of instances in long sequences to harness the embedded valuable information effectively. Image reconstruction is pivotal in enhancing diagnostic processes, as high-quality and high-fidelity medical images are fundamental to the accuracy and efficiency of clinical decisions. Huang et al.  [56] have developed MambaMIR, a model leveraging Mamba technology for the reconstruction of medical images, alongside its advanced counterpart, MambaMIR-GAN, which incorporates Generative Adversarial Networks. Moreover, Zheng et al. introduce FDVision Mamba (FDVM-Net)  [96], a frequency-domain-based network that effectively corrects image exposure by reconstructing the frequency domain of endoscopic images, as recorded endoscopic images often suffer from exposure abnormalities. In specialized areas, MD-Dose  [94], a cutting-edge diffusion model based on the Mamba architecture, was designed to predict radiation therapy dose distribution for thoracic cancer patients accurately.
@@
3D image analysis in medical imaging enables more accurate and comprehensive diagnoses by providing detailed views of complex anatomical structures. Gong et al. present nnMamba  [101], an innovative architecture designed for 3D medical imaging applications, which integrates local and global relationship modeling via the MICCSS (Mamba-In-Convolution with Channel-Spatial Siamese input) module. nnMamba was tested on a comprehensive benchmark of six datasets for three crucial tasks: segmentation, classification, and landmark detection, showcasing its capability in long-range relationship modeling at channel and spatial levels.